From today, how many months does Nour have to leave her money in the savings account for her to earn the $4700 she needs to buy a new machine for her cupcake shop?

Today, Nour’s savings are equal to $4100 and the savings account earns 6.50%.


1
Expert's answer
2021-01-25T11:54:21-0500

4700=4100(1+0.065)n

4700/4100 = 1.065n

introducing log

log(47/41) = n log 1.065

log(47/41)/Log 1.065 = n

n = 2.168 years

= 27 months